# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v1

@@ Abstract
Generating accurate SQL according to natural language questions (text-to-SQL) is a long-standing problem since it is challenging in user question understanding, database schema comprehension, and SQL generation. Conventional text-to-SQL systems include human engineering and deep neural networks. Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance. As modern databases become more complex and corresponding user questions more challenging, PLMs with limited comprehension capabilities can lead to incorrect SQL generation. This necessitates more sophisticated and tailored optimization methods, which, in turn, restricts the applications of PLM-based systems. Most recently, large language models (LLMs) have demonstrated significant abilities in natural language understanding as the model scale remains increasing. Therefore, integrating the LLM-based implementation can bring unique opportunities, challenges, and solutions to text-to-SQL research. In this survey, we present a comprehensive review of LLM-based text-to-SQL. Specifically, we propose a brief overview of the current challenges and the evolutionary process of text-to-SQL. Then, we provide a detailed introduction to the datasets and metrics designed to evaluate text-to-SQL systems. After that, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we discuss the remaining challenges in this field and propose expectations for future directions.
@@ I Introduction
Figure 1: An example for LLM-based text-to-SQL selected from Spider. The user proposes a question, “What cartoons were written by Joseph Kuhr?” the LLM takes the question and the schema of its corresponding database as the input, then generates a SQL query as the output. The SQL query can be executed in the database and retrieve a content “Batman Series” to answer the user question.
@@
Text-to-SQL is a long-standing task in natural language processing research. It aims to convert (translate) natural language questions into database-executable SQL queries. Fig. 1 provides an example of the large language model-based (LLM-based) text-to-SQL; given a user question “What cartoons were written by Joseph Kuhr?”, the LLMs take the question and its corresponding database schema as the input and generate an SQL query as the output, which can be executed in the database to retrieve the content “Batman Series” for answering the user question. The above system builds a natural language interface to the database (NLIDB) with LLMs. Since SQL remains one of the most widely used programming languages, with over half (51.52%) of professional developers using SQL in their work, but only around a third (35.29%) of those systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB enables non-skilled users to access structured databases like professional database engineers [1, 2] and also accelerates human-computer interaction [3]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can provide a solution to the prevalent hallucination [4, 5] issue by incorporating realistic content from the database to fill the knowledge gaps of LLMs [6]. The significant value and potential for text-to-SQL have triggered a range of studies on its incorporation and optimization with LLMs [7, 8, 9]; consequently, LLM-based text-to-SQL remains a highly discussed research field within the NLP and database communities.
@@
Previous studies have made notable progress in the implementation of text-to-SQL and have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [10, 11], specifically suitable for simple database scenarios. In recent years, with the heavy labor costs [12] brought by rule-based methods and the growing complexity of database environments [13, 14, 15], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [16, 17], which can automatically learn a mapping from the user question to its corresponding SQL [18, 19]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL [20], taking its performance to a new level [21, 22, 23]. Incrementally, PLM-based research focused on table content encoding [24, 19, 25] and pre-training [20, 11] has further advanced this field. Recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [7] and supervised fine-tuning (SFT) [9] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
@@
The overall implementation details of LLM-based text-to-SQL can be divided into 3 aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components in the database that match the user question; 3. SQL generation: This involves incorporating the above parsing and then writing correct syntax to generate executable SQL queries that can retrieve the desired answer. The LLMs have proven to perform a good vanilla implementation [26, 27], benefiting from the more powerful semantic parsing enabled by the richer training corpus compared to the PLMs [28, 29]. Further studies on enhancing the LLMs for question understanding [7, 8], schema comprehension [30, 31], and SQL generation [32] are being increasingly released.
@@
Figure 2: A sketch of the evolutionary process for text-to-SQL research from the perspective of implementation paradigm. Each stage is presented with some implementation techniques and representative works. The timestamps for the stages are not exactly accurate; we set each timestamp according to the release time of the representative works of each paradigm, with a margin of error of about one year before and after. The format is inspired from [29]
@@
Figure 3: Taxonomy tree of the research in LLM-based text-to-SQL. The display order in each node is organized by the released time. The format is adapted from [72]
I-A Challenges in Text-to-SQL
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [73]. In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art (SOTA) models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of LLMs and their application to text-to-SQL, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction to the recent advances of text-to-SQL incorporating LLMs. Specifically, the body of our survey covers a range of contents related to LLM-based text-to-SQL, including:
  • Datasets and Benchmarks: We provide an overview of the commonly used datasets and benchmarks for evaluating LLM-based Text-to-SQL systems. We discuss their characteristics, complexity, and the challenges they pose for text-to-SQL system development and evaluation.
  • Evaluation Metrics: We present the evaluation metrics used to assess the performance of LLM-based Text-to-SQL systems, including accuracy, exactness, and execution correctness. We discuss the advantages and limitations of each metric and their relevance to real-world applications.
  • Methods and Models: We explore the different methods and models employed for LLM-based text-to-SQL, including in-context learning and fine-tuning-based paradigms. We discuss their implementation details, strengths, and adaptations specific to the text-to-SQL task.
  • Expectations and Future Directions: We discuss the current challenges and limitations of LLM-based Text-to-SQL, such as real-world robustness, computational efficiency, data privacy and extensions. We also outline potential future research directions and opportunities for improvement.
We hope this survey will provide a clear overview of recent studies and inspire future studies. A taxonomy tree is shown in Fig. 3
@@ II Overview
Text-to-SQL is a task that aims to convert natural language-type questions into corresponding SQL queries that can be executed on a relational database. Formally, given a user question $\mathcal{Q}$ (also known as a user query, natural language, NL, NL question, etc.) and a database schema $\mathcal{D}$, the goal is to generate an SQL query $Y$ that can be executed on the database to obtain the desired answer. Text-to-SQL has the potential to democratize access to data by allowing users to interact with databases using natural language without the need for specialized knowledge of SQL. This can benefit various domains, such as business intelligence, customer support, and scientific research, by enabling non-technical users to easily retrieve information from databases and facilitating more efficient data analysis.
II-1 Linguistic Complexity and Ambiguity
@@
Natural language questions often contain complex linguistic structures, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to SQL queries. Additionally, natural language is inherently ambiguous, with multiple possible interpretations for a given question. Resolving these ambiguities and understanding the intent behind the question requires deep language understanding and the ability to incorporate context and domain knowledge.
II-2 Schema Understanding and Representation
@@
To generate accurate SQL queries, text-to-SQL systems need to have a comprehensive understanding of the database schema, including table names, column names, and relationships between tables. However, database schemas can be complex and vary significantly across different domains. Representing and encoding the schema information in a way that can be effectively utilized by the text-to-SQL model is a challenging task.
II-3 Rare and Complex SQL Operations
@@
Some SQL queries involve rare or complex operations, such as nested subqueries, outer joins, and window functions. These operations are less frequent in the training data and pose challenges for text-to-SQL models to generate accurately. Designing models that can handle a wide range of SQL operations, including rare and complex ones, is an important consideration.
II-4 Cross-Domain Generalization
@@
Text-to-SQL models often struggle to generalize across different database schemas and domains. Models trained on a specific domain may not perform well on questions from a different domain due to differences in vocabulary, schema structure, and question patterns. Developing models that can effectively adapt to new domains with minimal fine-tuning or domain-specific training data is an ongoing challenge.
II-A Evolutionary Process
@@
The field of text-to-SQL has witnessed significant advancements over the years, evolving from rule-based approaches to deep learning-based methods and, more recently, to the integration of pre-trained language models (PLMs) and large language models (LLMs), as shown in Fig. 2.
II-A1 Rule-based Approaches
@@
Early text-to-SQL systems relied heavily on rule-based approaches [10, 12, 11], where manually crafted rules and heuristics were used to map natural language questions to SQL queries. These approaches often involved extensive feature engineering and domain-specific knowledge. While rule-based methods achieved success in specific domains, they lacked the flexibility and generalization capabilities needed to handle diverse and complex questions.
@@
Post-annotated Dataset Release Time Base Dataset Special Setting Characteristics
@@
II-A2 Deep Learning-based Methods
With the rise of deep learning, sequence-to-sequence models, such as LSTMs and transformers, were adapted to generate SQL queries from natural language input [74, 19]. Typically, RYANSQL [19] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization. Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [18, 75].
II-A3 PLM-based Approaches
Pre-trained language models (PLMs) have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and understanding captured during pre-training. The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning pre-existing PLMs, such as BERT [24] and RoBERTa [76], on text-to-SQL datasets. These PLMs, which were pre-trained on large amounts of text data, captured rich semantic representations and language understanding capabilities. By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the knowledge and linguistic understanding of PLMs to generate accurate SQL queries [74, 20]. Another line of research focuses on incorporating schema information into PLMs to improve their understanding of database structures and generate more accurate SQL queries. Schema-aware PLMs are designed to capture the relationships and constraints present in the database schema [21].
II-A4 LLM-based Implementation
Large language models (LLMs), such as GPT series [77, 78], have gained significant attention in recent years due to their ability to generate coherent and fluent text. Researchers have started exploring the potential of LLMs for text-to-SQL by leveraging their vast knowledge and generative capabilities [26, 8]. These approaches often involve prompt engineering to guide the proprietary LLMs in the generation process or fine-tuning the open-source LLMs on text-to-SQL datasets.
The integration of LLMs in text-to-SQL is still an emerging area, and there is significant potential for further exploration and improvement. Researchers are investigating ways to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge, and develop more efficient fine-tuning strategies. As the field continues to evolve, we can expect to see more advanced and powerful LLM-based approaches that push the boundaries of text-to-SQL performance and generalization.
III Benchmarks & evalution
@@ III-A Datasets
As shown in Table I, we classify the datasets into ”Original Datasets” and ”Post annotated Datasets” based on whether they were released with the original dataset and databases or created by adapting existing datasets and databases with special processing. We also highlight their release times. For the original datasets, we provide a detailed analysis, including the number of examples, databases, tables per database, and rows per database. For the post-annotated datasets, we identify their base dataset and describe the special processing applied to them. To illustrate the potential opportunities of each dataset, we categorize their characteristics into Cross-domain, Knowledge-augmented, Context-dependent, Robustness, and Cross-lingual, which we will discuss in detail below.
III-A1 Cross-domain Dataset.
This refers to datasets where the background information of different databases comes from various domains. Since real-world text-to-SQL tasks often involve databases from multiple domains, most original text-to-SQL datasets [14, 13, 35, 34, 36, 33] and post annotated datasets [43, 42, 39, 40, 41, 37, 38] are in the cross-domain setting to fit well with the needs of cross-domain applications.

### v2

@@ Abstract
Generating accurate SQL according to natural language questions (text-to-SQL) is a long-standing challenge due to the complexities involved in user question understanding, database schema comprehension, and SQL generation. Conventional text-to-SQL systems, comprising human engineering and deep neural networks, have made substantial progress. Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance. As modern databases become more complex, the corresponding user questions also grow more challenging, leading PLMs with limited comprehension capabilities to produce incorrect SQL. This necessitates more sophisticated and tailored optimization methods for PLMs, which, in turn, restricts the applications of PLM-based systems. Most recently, large language models (LLMs) have demonstrated significant capabilities in natural language understanding as the model scale remains increasing. Therefore, integrating the LLM-based implementation can bring unique opportunities, improvements, and solutions to text-to-SQL research. In this survey, we present a comprehensive review of LLM-based text-to-SQL. Specifically, we propose a brief overview of the technical challenges and the evolutionary process of text-to-SQL. Then, we provide a detailed introduction to the datasets and metrics designed to evaluate text-to-SQL systems. After that, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we discuss the remaining challenges in this field and propose expectations for future research directions.
@@ I Introduction
Figure 1: An example for LLM-based text-to-SQL selected from the BIRD dataset. A user proposes a question about football leagues. The LLM takes the question and the schema of its corresponding database as the input and then generates an SQL query as the output. The SQL query can be executed in the database and retrieve the content “The 5 leagues with the highest matches” to answer the user question.
@@
Text-to-SQL is a long-standing task in natural language processing research. It aims to convert (translate) natural language questions into database-executable SQL queries. Fig. 1 provides an example of a large language model-based (LLM-based) text-to-SQL system. Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input and then generates an SQL query as output. This SQL query can be executed in the database to retrieve the relevant content to answer the user’s question. The above system builds a natural language interface to the database (NLIDB) with LLMs. Since SQL remains one of the most widely used programming languages, with over half (51.52%) of professional developers using SQL in their work, it is notable that only around a third (35.29%) of those developers are systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB enables non-skilled users to access structured databases like professional database engineers [1, 2] and also accelerates human-computer interaction [3]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can provide a potential solution to the prevalent hallucination [4, 5] issue by incorporating realistic content from the database to fill the knowledge gaps of LLMs [6]. The significant value and potential for text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [7, 8, 9, 10]; consequently, LLM-based text-to-SQL remains a highly discussed research field within the NLP and database communities.
@@
Previous studies have made notable progress in the implementation of text-to-SQL and have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [11], specifically suitable for simple database scenarios. In recent years, with the heavy labor costs [12] brought by rule-based methods and the growing complexity of database environments [13, 14, 15], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [16, 17], which can automatically learn a mapping from the user question to its corresponding SQL [18, 19]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL systems [20], boosting their performance to a new level [21, 22, 23]. Incremental research on PLM-based optimization, such as table content encoding [24, 19, 25] and pre-training [20, 26], has further advanced this field. Most recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [8] and fine-tuning (FT) [10] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
@@
The overall implementation details of LLM-based text-to-SQL can be divided into 3 aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components that match the user question; 3. SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer. The LLMs have proven to perform a good vanilla implementation [7, 27], benefiting from the more powerful semantic parsing capacity enabled by the richer training corpus [28, 29]. Further studies on enhancing the LLMs for question understanding [8, 9], schema comprehension [30, 31], and SQL generation [32] are being increasingly released.
@@
Figure 2: A sketch of the evolutionary process for text-to-SQL research from the perspective of implementation paradigm. Each stage is presented with two representative implementation techniques. The timestamps for the stages are not exactly accurate; we set each timestamp according to the release time of the representative works of each implementation paradigm, with a margin of error of about one year before and after. The format is inspired from [29].
@@
Figure 3: Taxonomy tree of the research in LLM-based text-to-SQL. The display order in each node is organized by the released time. The format is adapted from [72].
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [73]. In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of the implementation paradigm for text-to-SQL systems, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs. Specifically, the body of our survey covers a range of contents related to LLM-based text-to-SQL, including:
  • Datasets and Benchmarks: We provide a detailed introduction to the commonly used datasets and benchmarks for evaluating LLM-based text-to-SQL systems. We discuss their characteristics, complexity, and the challenges they pose for text-to-SQL development and evaluation.
  • Evaluation Metrics: We present the evaluation metrics used to assess the performance of LLM-based text-to-SQL systems, including both content matching-based and execution-based paradigms. We then briefly introduce the characteristics of each metric.
  • Methods and Models: We present a systematic analysis of the different methods and models employed for LLM-based text-to-SQL, including in-context learning and fine-tuning-based paradigms. We discuss their implementation details, strengths, and adaptations specific to the text-to-SQL task from various implementation perspectives.
  • Expectations and Future Directions: We discuss the remaining challenges and limitations of LLM-based text-to-SQL, such as real-world robustness, computational efficiency, data privacy, and extensions. We also outline potential future research directions and opportunities for improvement and optimization.
We hope this survey provides a clear overview of recent studies and inspires future research. Fig. 3 shows a taxonomy tree that summarizes the structure and contents of our survey.
@@ II Overview
Text-to-SQL is a task that aims to convert natural language questions into corresponding SQL queries that can be executed in a relational database. Formally, given a user question $\mathcal{Q}$ (also known as a user query, natural language question, etc.) and a database schema $\mathcal{S}$, the goal of the task is to generate an SQL query $Y$ that retrieves the required content from the database to answer the user question. Text-to-SQL has the potential to democratize access to data by allowing users to interact with databases using natural language without the need for specialized knowledge of SQL programming [74]. This can benefit various domains, such as business intelligence, customer support, and scientific research, by enabling non-skilled users to easily retrieve target content from databases and facilitating more efficient data analysis.
II-A Challenges in Text-to-SQL
The technical challenges for text-to-SQL implementations can be summarized as follows:
II-A1 Linguistic Complexity and Ambiguity
@@
Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding part of SQL queries [41]. Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [75, 76]. Resolving these ambiguities and understanding the intent behind the user question requires deep natural language understanding and the capability to incorporate context and domain knowledge [33].
II-A2 Schema Understanding and Representation
@@
To generate accurate SQL queries, text-to-SQL systems need to have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables. However, the database schema can be complex and vary significantly across different domains [13]. Representing and encoding the schema information in a way that can be effectively utilized by the text-to-SQL model is a challenging task.
II-A3 Rare and Complex SQL Operations
@@
Some SQL queries involve rare or complex operations and syntax in challenging scenarios, such as nested sub-queries, outer joins, and window functions. These operations are less frequent in the training data and pose challenges for text-to-SQL systems to generate accurately. Designing models that can be generalized to a wide range of SQL operations, including rare and complex scenarios, is an essential consideration.
II-A4 Cross-Domain Generalization
@@
Text-to-SQL systems often struggle to generalize across various database scenarios and domains. Models trained on a specific domain may not perform well on the proposed questions from other domains due to the variety in vocabulary, database schema structure, and question patterns. Developing systems that can effectively generalized to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [77].
II-B Evolutionary Process
@@
The research field of text-to-SQL has witnessed significant advancements over the years in the NLP community, having evolved from rule-based methods to deep learning-based approaches and, more recently, to integrating pre-trained language models (PLMs) and large language models (LLMs), a sketch of the evolutionary process is shown in Fig. 2.
II-B1 Rule-based Methods
@@
Early text-to-SQL systems relied heavily on rule-based methods [11, 12, 26], where manually crafted rules and heuristics were used to map natural language questions to SQL queries. These approaches often involved extensive feature engineering and domain-specific knowledge. While rule-based methods achieved success in specific simple domains, they lacked the flexibility and generalization capabilities needed to handle diverse and complex questions.
II-B2 Deep Learning-based Approaches
With the rise of deep neural networks, sequence-to-sequence models and encoder-decoder structures, such as LSTMs [78] and transformers [17], were adapted to generate SQL queries from natural language input [79, 19]. Typically, RYANSQL [19] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization. Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [18, 80].
II-B3 PLM-based Implementation
Pre-trained language models (PLMs) have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and semantic understanding captured during the pre-training process. The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning off-the-shelf PLMs, such as BERT [24] and RoBERTa [81], on standard text-to-SQL datasets [13, 14]. These PLMs, pre-trained on large amounts of training corpus, captured rich semantic representations and language understanding capabilities. By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the semantic and linguistic understanding of PLMs to generate accurate SQL queries [79, 20, 82]. Another line of research focuses on incorporating schema information into PLMs to improve their understanding of database structures and enable them to generate more executable SQL queries. Schema-aware PLMs are designed to capture the relationships and constraints present in the database structure [21].
II-B4 LLM-based Implementation
Large language models (LLMs), such as GPT series [83, 84, 85], have gained significant attention in recent years due to their capability to generate coherent and fluent text. Researchers have started exploring the potential of LLMs for text-to-SQL by leveraging their extensive knowledge reserves and superior generation capabilities [7, 9]. These approaches often involve prompt engineering to guide the proprietary LLMs in SQL generation [47] or fine-tuning the open-source LLMs on text-to-SQL datasets [9].
The integration of LLMs in text-to-SQL is still an emerging research area with significant potential for further exploration and improvement. Researchers are investigating approaches to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge [33, 31], and develop more efficient fine-tuning strategies [10]. As the field continues to evolve, we anticipate the development of more advanced and superior LLM-based implementations that will elevate the performance and generalization capabilities of text-to-SQL to new heights.
III Benchmarks & evalution
@@
Post-annotated Dataset Release Time Source Dataset Special Setting Characteristics
@@ III-A Datasets
As shown in Tab. I, we classify the datasets into ’Original Datasets’ and ’Post-annotated Datasets’ based on whether they were released with the original dataset and databases or created by adapting existing datasets and databases with special settings. For the original datasets, we provide a detailed analysis, including the number of examples, the number of databases, the number of tables per database, and the number of rows per database. For the post-annotated datasets, we identify their source dataset and describe the special setting applied to them. To illustrate the potential opportunities of each dataset, we annotate them based on their characteristics. The annotations are listed in the rightmost column of Tab. I, which we will discuss in detail below.
III-A1 Cross-domain Dataset
This refers to datasets where the background information of different databases comes from various domains. Since real-world text-to-SQL applications often involve databases from multiple domains, most original text-to-SQL datasets [14, 13, 35, 34, 36, 33] and post-annotated datasets [43, 42, 39, 40, 41, 37, 38] are in the cross-domain setting to fit well with the requirements of cross-domain applications.